OpenStudy (anonymous):

What's the difference between geometric and arithmetic sequences? How do you put them in summation notation?

OpenStudy (amistre64):

what definitions do they give you?

OpenStudy (anonymous):

We don't have any I just don't know what they are in general

OpenStudy (amistre64):

arithmetic is adding ... you are adding the same value one term to get to the next term geometric is similar, but its multiplying the value instead of adding

OpenStudy (amistre64):

a summation notation is simply taking the rule that defines the sequence, and adding all the terms that the sequence generates

OpenStudy (amistre64):

for example: lets add 3 to some seed value: 2 2 2+3 = 5 5+3 = 8 8+3 = 11 11+3 = 14 etc ....

OpenStudy (anonymous):

so squence would be 2, 5, 8, 11, 14 a1, a2, a3, a4, a5

OpenStudy (amistre64):

yes, sometimes its convenient to start at n=0, but thats not important :)

OpenStudy (anonymous):

maybe i'll actually pass my math quiz

OpenStudy (amistre64):

notice that we can make a rule for the sequence. a1 = 2 + 3(0) a2 = 2 + 3(1) a3 = 2 + 3(2) a4 = 2 + 3(3) a5 = 2 + 3(4) for every value of n, we add an (n-1) multiple to our seed value

OpenStudy (amistre64):

so the rule that explicitly defines this sequence that we are generating is: 2 + 3(n-1)

OpenStudy (amistre64):

lets apply this concept to a geometric sequence: a common ratio is the conventioanl term to use for the common value. lets call it 'r' with the same seed value of 2 g1 = 2r^0 g2 = 2r^1 g3 = 2r^2 g4 = 2r^3 g5 = 2r^4 in general, the explicit rule for this geometric sequence is: 2r^(n-1)

OpenStudy (amistre64):

any questions?

OpenStudy (anonymous):

I think I got it all I guess we'll see how we do

OpenStudy (amistre64):

goodluck ;)